A car drops off a cliff that has a height of 700m. How long does it take for the car to hit the ground?
1
Expert's answer
2020-11-30T14:53:34-0500

We can find the time that the car takes to hit the ground from the kinematic equation:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot 700\ m}{9.8\ \dfrac{m}{s^2}}}=11.95\ s."

Answer:

"t=11.95\ s."